40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Mon Nov 30 11:28:30 2020, aaron, Noise Hunting, Laser, pll IMG_0195.jpg
    Reply  Tue Dec 1 10:59:49 2020, aaron, Noise Hunting, Laser, pll 
       Reply  Wed Dec 2 13:37:55 2020, aaron, Noise Hunting, Laser, W laser oscillating 
          Reply  Thu Dec 3 10:18:07 2020, aaron, Noise Hunting, Laser, pll 
          Reply  Thu Dec 3 11:57:52 2020, rana, Noise Hunting, Laser, W laser oscillating 
             Reply  Fri Dec 4 10:36:18 2020, aaron, Noise Hunting, Laser, pll IMG_0015.PNG
                Reply  Sun Dec 6 19:22:10 2020, rana, Noise Hunting, Laser, pll 
                   Reply  Thu Dec 10 10:05:03 2020, aaron, Noise Hunting, Laser, pll IMG_0016.PNG
       Reply  Wed Feb 10 12:39:08 2021, rana, Noise Hunting, Laser, pll 
          Reply  Thu Feb 11 09:59:31 2021, aaron, Noise Hunting, Laser, pll 
          Reply  Thu Feb 11 17:24:41 2021, rana, Noise Hunting, Laser, pll PXL_20210212_000327036~2.jpg
             Reply  Mon Feb 15 14:57:25 2021, aaron, Noise Hunting, Laser, pll 
Message ID: 2596     Entry time: Wed Dec 2 13:37:55 2020     In reply to: 2595     Reply to this: 2597   2598
Author: aaron 
Type: Noise Hunting 
Category: Laser 
Subject: W laser oscillating 

Enter Wed Dec 2 13:37:57 2020

I'd like to know whether the problem is in the W laser driver (S1500207) or the diode (104987). I swap the E laser drive over to the W laser diode, so am driving diode 104987 with driver S1600246. I still don't see the E laser driver oscillating, but the W diode power still reaches only 1.38 mW before falling off (ramping up current at T_set). Just in case, I also drive the W diode using the E driver but the cable from the original W driver (checking that the cable isn't the cause); no oscillation.

Entered EE to obtain grounding strap (Wed Dec 2 14:10:12 2020). On second thought, putting off diagnosing the W laser drive and just continuing with ITC 502. I did get a look inside, and don't actually see evidence that R33 was changed, just a added resistor from R30 to V_ref mentioned on the schematic.

Found this elog helpful for making sure the pinout for ITC 502 is compatible. Oddly, I can't set the laser diode current limit higher than 10mA, though the range of the device is +- 200 mA. I suspect the left 2 digits of the current display are just inoperable, because I can turn the adjustment pot in either direction until the digits stop moving. I set the current limit at 149 mA by counting by 10, and will double check the eventual current from the monitor. the diode definitely is nearing end of life, I'm only getting 1mW at nominal setpoint. I was squeezing a little extra out of the custom laser driver, but probably at the expense of the diode.

With the thorlabs driver, we're down to 1 peak at last!

Looking at the spectrum today, maybe the low frequency cutoff I noticed yesterday is just a frequency offset issue?

switching over to Moku... wow this is a user friendly device.

  • beat note drifts by MHz over minute timescales
  • FWHM ~ 140 kHz
  • 35 dB above noise floor

Phaselock setup

  • Photodiode output to splitter (ZFSC-2-4+)
    • Long cable to 83 dB (!!?) gain amplifier to mixer (ZAY-1B) LO input. Mixer requires +23 dBm LO.
    • short cable to [gain] to mixer (ZAY-1B) RF input

While choosing the gain for producing an LO from the 1611 output, I tried to measure the absolute power with the Moku. When I change the range from 1 Vpp to 10 Vpp, the overall peak height changes by almost 10 dB. I wondered if this were just a clipping issue, so I hooked up the Moku function generator output to the channel 2 input to see if the behavior is consistent. The beat note is 300 mVpp at the laser settings I'm currently using, which should be about 0 dBm dissipated into 1 MOhm. Sourcing 300 mVpp from the Moku into its channel 2 input gives the correct power reading, but the 300 mVpp from the 1611 reads -25 dBm with 10 Vpp range and -34 dBm with 1 Vpp range. What's going on?

  • Probably just outside the linear range of the 1611, which saturates at 1mW input power. I'm seeing some unexpected sidebands (>> 9MHz), so nonlinearity makes sense. We're still well below the 10 mW damage threshold, but should add back the 90-10 pickoff mirrors now that alignment is done.
  • After adding the 90-10 (power) BS, the AC output of the 1611 has ~3 mVpp (amplitude). nice
  • The beat is now -53 dBm. That's still below what I'd expect for a 3 mVpp signal; Moku thinks it's seeing a 1.45 mVpp
    • I want to attribute this to some setting of the oscilloscope or Moku, but it's also not consistent with my expectation. moku shows 6 mVpp when measuring its 3 mVpp output with 1 MOhm input impedance, suggesting there's a factor of 2 I'm missing somewhere; on the other hand, the signal from the 1611 needs a factor of 1/2 to be consistent with the oscilloscope measurement. Awkward.
  • started these measurements at 50 Ohm input impedance for the SA, but when results didn't make sense I repeated at 1 MOhm. Should have stuck with 50.

I would have to string together eg 2x SR560 to get enough gain. I guess that's fine, because the noise floor of the SR560 is O(nV/rtHz)... but it gives me the feeling of doing something wrong, and I'm hungry so will pick it up tomorrow.

exit Wed Dec 2 18:42:15 2020

ELOG V3.1.3-